# Distributed Multisensor ISAC

### v1

@@ V-A Multistatic Dynamic Target State Vector Estimation
The underlying spatially distributed localization paradigm has inherent advantages over the single station range/bearing localization of monostatic ISAC sensors. Single station monostatic radars suffer from decreasing SNR and corresponding probability of detection by the fourth power of the radial distance according to the radar equation. In addition, the cross range resolution decreases with increasing distance from the target. While increasing transmit power beyond the regulatory limit is not an appropriate approach for ISAC, more degrees of freedom to balance the spatial distribution of performance arise with distributed ISAC. The spatial coverage of MS ISAC becomes evident from the Cassini ovals, which describe the coverage maps of bistatic radars. The $0\text{\,}\mathrm{dB}\text{/}$ level in Figure 9 refers to direct transmission from Tx to Rx and thus to the Friis equation. Therefore, the decibels describe the additional attenuation due to excessive range. We deduce that the best SNR is achieved when the target is close to one ISAC radio node (Tx or Rx). It is moderate when the target is located somewhere between the two nodes and it decreases steeply according to the radar equation when it is far away from both nodes. Therefore, an advantageous sensor scenario with drones could be that the drone approaches the target while the illuminator is a powerful gNB in its fixed position.
@@
As explained in the previous sections, MS ISAC performance results from cooperation of multiple distributed radio nodes, respectively from fusion of their measurements. Therefore, the most obvious approach for enhancing and smoothing area-wide coverage and sensing performance is network densification, as we already know from mobile communications. By choosing the relative position of the ISAC nodes, we can achieve a smooth coverage and resolution performance by controlling of the SNR and GDOP parameters. We can even adjust the performance according to the local needs. Heterogeneous nodes can help further. Although multilateration can be seen as an advantage if we do not have a DoA estimate available, beamforming can be used to increase SNR and, hence, improve performance in certain situations. Different frequency bands can be used, e.g. to achieve higher performance at target hot spots with nodes using millimeter wave frequencies or to guarantee minimum sensing performance in less critical situations by auxiliary umbrella nodes at lower frequencies.
@@
Last but not least, distributed ISAC is better suited for 3D target localization as it has the potential to capture all spatial coordinates. For instance, although single station range / bearing radar can act as an autonomous localization system, it is “Doppler blind” for cross range moving targets.
@@
Figure 8: Two crossing ellipses with small and large GDOP. Wheen the ellipses intersect at a right angle (Tx,Rx1 & Tx,Rx3), the position error is small (yellow); where they intersect at a flat angle (Tx,Rx1 & Tx,Rx2), it is large (violet).
@@ V-B MS Radio Access and Scheduling
Obviously, MS-MIMO radio access needs coordination and scheduling. The multiple radio links in Figure 5 share the same radio environment. Therefore we need orthogonal medium access schemes to make sure that the multiple links do not create mutual interference. The links should permit multiple (quasi) simultaneous measurements because of the dynamic 3D geometric nature of the problem. These measurements undergo bandwidth and real-time constraints and require adequate temporary coordination because of the dynamic evolution of the whole scenario. Requirements on range-Doppler resolution and coherent radar integration time places demands on bandwidth, synchronization, and sustained recording time. Eventually, multi-antenna access focused in beam domain and matched in polarization at Tx and/or Rx side would be necessary. With the limited resources of the shared medium, we are therefore subject to fundamental performance limits that would indicate how many dynamic targets in a given environment we can reliably detect, resolve and track. Finally yet importantly, in addition to the sensing performance of the physical radio interface, we may have to find a performance balance of ISAC with communication services about radio and other network resources, which also includes demands for sensing data transport and fusion.
@@
While some basic aspects of radio access have already been discussed in the previous chapter on architectural issues, it becomes clear that we need to differentiate MS-MIMO access into MS broadcast and orthogonal MS access. MS broadcast access (preferably in DL), where only one radio node in Figure 5 is transmitting and the others are listening, provides several measurements at the same time, which is very advantageous for dynamic targets although it does not cover all possible sensing links (only N1N-1 out of $N^{2}$), see Figure 10. The orthogonal MS mode turns the broadcast MS channel into a shared sensing channel. We have to rely on nested distribution of Resource Elements (REs), also referred to as Resource Blocks (RBs) in the time-frequency plane (OFDMA in DL or SC-FDMA in UL). The orthogonal MS mode turns the broadcast MS channel into a shared sensing channel.
@@
Another problem is related to the TDD scheme in 5G NR. Dynamic TDD DL/UL-split is aimed at supporting assignments based on instantaneous traffic demands. Consequently, there is a gap in one TDD stream when the other is active, and normally the duty cycle of the DL is greater than that of the UL. This raises two questions. The first is which influence do the RE distribution and the DL/UL balance have for the dynamic target estimation performance and what is the optimum allocation of RE and TDD gaps? The second results in an estimation problem from sparse data. Sparse allocation of RE and TDD gaps renders simple 2D-FFT target estimators useless since replacing the missing samples in frequency domain just by zeros would distort the desired shape of the Ambiguity Function (AF) of the correlation reference waveform by disturbing sidelobes.
@@
The flexibility of 5G New Radio already allows for great waveform adaptability and optimum coordination of ISAC radio access. In addition to the physical OFDM parameters defined by the numerology, it is about the time-frequency distribution of the REs and TDD gaps in the radio frame which is crucial for resource allocation. This enables the 5G communication system to accommodate a variety of applications with different bandwidth, latency, and reliability needs. We will observe similar influence of time-frequency resource distribution for radar sensing quality control. ISAC radar sensing quality parameters are related to sensitivity of target detection, target resolution in range and Doppler, robustness against misdetection in clutter, target classification and tracking performance in dynamic scenarios
@@
Here we will discuss some basic influential factors of MS access schemes and MS RE distribution. Figure 11 depicts the same three-node scenario as in Figure 10, however with two transmitters and one receiver. The two transmitters share the same Orthogonal Frequency-Division Multiple Access (OFDMA) frame. We will recognize that sharing these resources also means a certain division of resolution performance for range and/or Doppler. The optimum sharing of OFDMA resources depends on the geometry of the sensor cluster relative to the target. In the left case, the direction of the targets velocity vector is tangential to the Tx1/Rx ellipse while the right case aligns with the Tx2/Rx ellipse. Provided the radio nodes are stationary, the observed Doppler shift would be zero for the Tx1/Rx link in the left case and zero for the Tx2/Rx in the right case. At the same time, in the corresponding other cases the observed Doppler shift would be higher, although not necessarily maximum. The maximum Doppler shift would be observed, when the velocity vector is perpendicular to the ellipses. This situation can be exploited at the transmitters for resource block allocation. Figure 12 depicts two examples of slow time sampling by OFDMA resource blocks with four transmitters. The left case describes uniform, repetitive sampling at the same rate for all four Tx. We see that shared TDMA access to the medium reduces the sampling rate in slow time and thus the maximum Doppler shift that can be estimated unambiguously. The included timing offset can be compensated for as long as the Nyquist sampling rate is maintained. The right part of Figure 12 explains how the TDMA scheme can be adjusted to the target dynamics. While Tx3 illuminates the target still with the same sampling rate in slow time, the unambiguously resolved Doppler bandwidth has reduced for Tx2 and Tx4. Tx1 can be used in two ways. If the two neighboring OFDM symbols are averaged, the effective overall Doppler sampling rate is kept, the Doppler bandwidth is reduced only a bit and the sensitivity of target detection increased. If the consecutive pairs of OFDM symbols undergoes the FFT, the SNR gain by coherent averaging still applies, but we will also have some possibility to increase resolvable Doppler frequency. The sensitivity of target detection can also be influenced by the radar integration time, i.e. the length of the slow time observation window, which can extend over several coherently processed OFDM frames. The limit is given by range migration, if the target moves from one range cell to the next one (in range-Doppler domain).
@@
Figure 13 explains the FDMA split within the OFDM frame. The left part shows a uniform bandwidth allocation with compact bands that enable uniform range resolution for all four sensor links. In the right part, the same total bandwidth is allocated per transmitter, but the allocation is not consecutive, except for Tx4. A closer look shows that the range resolution capability for Tx1 is highest. However, because of the fragmented (”sparse”) frequency allocation straightforward FFT estimation would not produce viable estimates because of grating lobes in the autocorrelation function. We would need a more sophisticated estimation procedure than just FFT that can cope with the sparse allocation of frequency resources as described in more detail in the next subsection.
@@
Some further comments are necessary. The time-frequency allocation described here as an example does not consider the time-frequency selectivity of the sensing channel. Especially the frequency selectivity because of multipath can require a more specific allocation of REs. Moreover, adaptive REs scheduling for joint optimization of sensing and communication quality of service would be necessary. Another interesting observation comes from the fact that for localization we would need multiple measurements, e.g. at least 3 Tx/Rx pairs in for 3D location and speed estimation. This means that the optimal distribution of resource blocks in the OFDM frame depends on the sensor geometry relative to the target and is mutually dependent. Moreover, for multiple targets it depends on the difference of their dynamic target state vectors in its multiple dimensions. Therefore, the optimal allocation of resources in the OFDM frame leads to a joint optimization problem depending on the positions of all sensors used and the dynamic behavior of all relevant targets given the available resources.
@@
Figure 11: Orthogonal MS access with Tx1 and Tx2 sharing the same OFDMA frame. In both cases the position of the target is the same, but the direction of its movement is different.
@@
The discussion of MS-MIMO access in ISAC and any comparison with MU MIMO communication would be incomplete without considering Joint Transmission (JT), where the same illumination signal is transmitted by multiple distributed transmitters. Already throughout research towards 4G, cooperative multipoint access (CoMP) including JT was discussed, especially for enhanced quality of service at cell edges. Although successfully demonstrated [ref33], it seems that CoMP has not been widely used up to now. However, the situation may develop differently for ISAC because of its inherent multilink requirements as explained above. JT for MS-MIMO ISAC definitely deserves attention in research. There seem to be remarkable differences to JT in communications. We have to make sure that focusing only involves the target and not the clutter. This could be achieved by background subtraction, e.g. if the target is moving and the clutter is static. However, due to the statistical reflectivity of typical extended targets, which leads to decorrelated radar returns as described above, coherent focusing to certain reflection points would only be possible for single Tx/Rx pairs. Therefore, although coherent JT promises tremendous spatial resolution that could resolve details from extended targets down to single reflection points, non-coherent focusing seems to be a more reasonable goal. This would lead to increased reflected energy off the target without increasing spatial resolution. At the same time, we get a diversity gain because of the different directions of the waves impinging to the target [ref34]. Thus, non-coherent JT ISAC seems to keep the statistical advantage of distributed MIMO radar, but the geometrical advantage may get lost since the multistatic Tx/Rx geometry cannot be traced back if the non-orthogonal waveforms cannot be decomposed at the receiver. However, since the Channel State Information (CSI) pilots are orthogonal, simultaneous estimation of the joint target channels is possible which would allow adjusting the transmit signal delay in a subsequent step for non-coherent superposition as indicated in Figure 14.
@@ V-C Recovery of the Correlation Reference Signal
Excess time of arrival estimation, also jointly together with Doppler, can be based on specific pilot signals or on using the data payload as correlation reference to compare LOS path and BTP time of flight. Although pilots and other reference signals are designed for estimation of CSI in communications or UE positioning, they may be not enough for high dynamic range target state vector estimation in high dynamic range radar sensing. One reason for this is their sparse structure, which is sufficient for equalization, but interpolation errors in frequency domain and the related distortion of the AF may not be tolerated for radar sensing. Moreover, since CSI pilots use only a fraction of the transmit signal power the achievable correlation gain is limited. On the other hand, asking for more pilot signals is counterproductive, since this would reduce communication efficiency too much. Even extensive use of Position Reference Signals (PRS) would have the same effect. In contrast to this, CPCL as proposed in [ref1] and introduced already here in Section II, relies on the full data payload for precise target estimation and tracking. Note that radar localization has a systematic disadvantage compared to 3GPP UE device positioning since it relies on passive backscattering. This means that the sensing link budget follows the radar equation and not the Friis equation. Furthermore, the target cannot be identified by some user ID. Instead, we need to use much more sophisticated means to detect, recognize and separate multiple targets. Using the data payload as correlation reference offers not only the possibility to use the total transmitted power (maximizing the SNR) and system bandwidth (maximizing resolution). It also allows the correlation reference waveform to be adjusted and modified, enabling unprecedented approaches to adaptive radar functions. However, the transmitted payload is not known a priori at the receiver (unlike the pilots) and has to be estimated at the receiver before performing correlation. While conventional passive radar (PCL) uses a dedicated, directional receiver channel to obtain a precise, multipath-free copy of the transmitted signal via LOS, ISAC’s own communication function offers many more options for recovering the transmitted signal without a separate Rx channel. However, we have to be careful in defining what we need as a reference and which recovering approach we apply. We will discuss different approaches and how they build upon the standard OFDM transmission approach.
@@
In addition to Tx/Rx frame and carrier synchronization and cyclic prefix removal, an OFDM receiver performs channel equalization infrequency domain, which provides a symbol-by-symbol estimate of the transmitted OFDM signal based on multiplication by the inverse channel frequency response estimated at the CSI pilot positions and then interpolated to apply to all carriers. This procedure is followed by decision making out of a finite modulation alphabet and by error correcting channel coding. As a result, the transmitted information can be decoded at the receiver with very low bit error probability. Therefore, we might come up with the idea to use that information for recovering the transmitted waveform for ISAC correlation by remodulation. However, the requirements for reliable and high dynamic range radar target detection and estimation is a bit different from communications. While in communication we detect coded information per carrier, in radar we perform a matched filter correlation using the whole OFDM symbol to identify target range. The OFDM symbol is treated as an extended spread spectrum radar pulse that is compressed by correlation, which inheres a rather high correlation gain. Moreover, we take further advantage from coherent processing of many subsequent OFDM symbols. The result should clearly indicate the target related propagation parameters delay and Doppler as described in the next subsection. However, the target maybe associated only with a very small part of the received energy, perhaps in the neighborhood of strong clutter peaks and hidden by their correlation sidelobes. The dynamic range is therefore generally higher for ISAC than for communication. High detection probability and good resolution of target parameters in a continuous parameter space depends more on precise knowledge of the illumination waveform and its coherency over time than on its information content in terms of bits and bytes. However, RF impairments at the transmitter can affect the waveform and, hence, the shape of its correlation function. For instance, since with OFDM the power amplifier is often operated close to saturation, non-linear distortions can occur [ref37]. Also linear distortions (complex frequency response, IQ imbalance, etc.) of the modulator, transmit filters, and antennas can prove to be disruptive. For communications, those effects are normally leveled out by the inherent error correction mechanisms. This may not be true for radar, as the detection problem is different. Targets often appear as small peaks in environments with strong clutter. They only become apparent when overlapping system responses can be resolved through iterative subtraction using a model-based estimator, as described in the next subsection. Therefore, precise knowledge of the shape of the transmitted waveform, as well as its correlation function (including the sidelobes), is of the utmost importance. This includes all distortions of the RF chain, including nonlinear ones that appear after digital reconstruction. Thus, it is preferable to rely entirely on the recovery of the correlation reference received over the air for target parameter estimation.
@@
Inverse filtering is the most important step in correlation reference recovery because it equalizes the frequency response, removes associated multipath and provides a clean copy of the transmitted signal. However, the usual inverse filter operation in the frequency domain has serious pitfalls. One relates to REs (REs p. carriers) which are loaded with zero power. First of all, these empty REs should not be included in the inverse filter operation. In the fully connected CPCL mode, the information about empty REs can be directly taken from the DeModulation Reference Signal (DMRS). In the more general case of the reduced CPCL mode, empty REs in the received reference signal are detected at the receiver based on a null hypothesis test [ref38]. Another related issue when recovering the transmit signal at the receiver side is the non-uniform (instantaneous) carrier magnitude. A simple inverse filter equalizes all received carriers to unit gain of the channel transfer function. This would lead to a non-uniform noise distribution of the carriers used for target parameter estimation. Since the CSI pilots are usually provided with more power and more robust modulation than the data carriers, these estimates would have a lower variance. On the other hand, multilevel QAM carrier modulation would lead to noise amplification at the carrier positions that are loaded according to the modulation alphabet by a lower magnitude of the constellation diagram. These effects has to be taken into account for optimum target parameter estimation as discussed in the next subsection.
@@
The quality of the reconstructed transmit signal can be further enhanced by a two step iterative procedure. In a first step, we perform regular equalization of the received REs based on CSI pilots. The result is used for a first inverse filter step as described above. While these estimated payload data are processed for communication data reception, we use them to perform another channel equalization that should now be more accurate as the regular equalization since it uses the full transmit power and should be less affected by interpolation errors, which can arise because of the sparse structure of CSI pilots. With the now enhanced reference signal, we can calculate the final channel impulse response, which is used for radar target detection and estimation. The advantage from this approach compared to the alternative based on transmit symbol decision, decoding, and remodulation in the receiver is as follows. All linear and non-linear distortions at the transmitter are now included in the correlation reference which assures a high dynamic range estimate. Even the time reference is kept and Tx2Rx synchronization is automatically established since the critical eToA and eDoppler are effectively calculated as differences between LOS and BTP from the same measured CIR (see discussion on synchronization issues later on). We refer to this approach as Turbo CPCL because it is a nested iterative approach that takes advantage of the processing gain in the inner estimation loop using the CSI pilots. In terms of estimation accuracy, we can also understand the outer loop as a refinement of the first estimate.

### v2

@@ V-A Multistatic Dynamic Target State Vector Estimation
The underlying spatially distributed localization paradigm has inherent advantages over the single station range/bearing localization of monostatic ISAC sensors. Single station monostatic radars suffer from decreasing SNR and corresponding probability of detection by the fourth power of the radial distance according to the radar equation. In addition, the cross-range resolution decreases with increasing distance from the target. While increasing transmit power beyond the regulatory limit is not an appropriate approach for ISAC, more degrees of freedom to balance the spatial distribution of performance arise with distributed ISAC. The spatial coverage of MS-ISAC becomes evident from the Cassini ovals, which describe the coverage maps of bistatic radars. The $0\text{\,}\mathrm{dB}\text{/}$ level in Figure 9 refers to direct transmission from Tx to Rx and thus to the Friis equation. Therefore, the decibels describe the additional attenuation due to excessive range. We deduce that the best SNR is achieved when the target is close to one ISAC radio node (Tx or Rx). It is moderate when the target is located somewhere between the two nodes, and it decreases steeply according to the radar equation when it is far away from both nodes. Therefore, an advantageous sensor scenario with drones could be that the drone approaches the target while the illuminator is a powerful gNB in its fixed position.
@@
As explained in the previous sections, MS-ISAC performance results from the cooperation of multiple distributed radio nodes, respectively, from the fusion of their measurements. Therefore, the most obvious approach for enhancing and smoothing area-wide coverage and sensing performance is network densification, as we already know from mobile communications. By choosing the relative position of the ISAC nodes, we can achieve a smooth coverage and resolution performance by controlling the SNR and GDoP parameters. We can even adjust the performance according to the local needs. Heterogeneous nodes can help further. Although multilateration can be seen as an advantage if we do not have a DoA estimate available, beamforming can be used to increase SNR and, hence, improve performance in certain situations. Different frequency bands can be used, e.g., to achieve higher performance at target hot spots with nodes using millimeter wave frequencies or to guarantee minimum sensing performance in less critical situations by auxiliary umbrella nodes at lower frequencies.
@@
Last but not least, distributed ISAC is better suited for 3D target localization as it has the potential to capture all spatial coordinates. For instance, although a single station range / bearing radar can act as an autonomous localization system, it is “Doppler blind” for cross-range moving targets.
@@
Figure 8: Two crossing ellipses with small and large GDoP. When the ellipses intersect at a right angle (Tx,Rx1 & Tx,Rx3), the position error is small (yellow); where they intersect at a flat angle (Tx,Rx1 & Tx,Rx2), it is large (violet).
@@ V-B MS Radio Access and Scheduling
Obviously, MS-MIMO radio access needs coordination and scheduling. The multiple radio links in Figure 5 share the same radio environment. Therefore, we need orthogonal medium access schemes to make sure that the multiple links do not create mutual interference. The links should permit multiple (quasi) simultaneous measurements because of the dynamic 3D geometric nature of the problem. These measurements undergo bandwidth and real-time constraints and require adequate temporary coordination because of the dynamic evolution of the whole scenario. Requirements on range-Doppler resolution and coherent radar integration time place demands on bandwidth, synchronization, and sustained recording time. Eventually, multi-antenna access focused in the beam domain and matched in polarization at Tx and/or Rx side would be necessary. With the limited resources of the shared medium, we are therefore subject to fundamental performance limits that would indicate how many dynamic targets in a given environment we can reliably detect, resolve, and track. Finally, yet importantly, in addition to the sensing performance of the physical radio interface, we may have to find a performance balance of ISAC with communication services about radio and other network resources, which also includes demands for sensing data transport and fusion.
@@
While some basic aspects of radio access have already been discussed in the previous chapter on architectural issues, it becomes clear that we need to differentiate MS-MIMO access into MS broadcast and orthogonal MS access. MS broadcast access (preferably in DL), where only one radio node in Figure 5 is transmitting and the others are listening, provides several measurements at the same time, which is very advantageous for dynamic targets although it does not cover all possible sensing links (only N1N-1 out of $N^{2}$), see Figure 10. The orthogonal MS mode turns the broadcast MS channel into a shared sensing channel. We have to rely on the nested distribution of Resource Elements (REs), also referred to as Resource Blocks (RBs) in the time-frequency plane (Orthogonal Frequency-Division Multiple Access (OFDMA) in DL or SC-Frequency-Division Multiple Access (FDMA) in UL). The orthogonal MS mode turns the broadcast MS channel into a shared sensing channel.
@@
Another problem is related to the TDD scheme in 5G NR. Dynamic TDD DL  / UL-split is aimed at supporting assignments based on instantaneous traffic demands. Consequently, there is a gap in one TDD stream when the other is active, and normally the duty cycle of the DL is greater than that of the UL. This raises two questions. The first is which influence do the RE distribution and the DL  / UL balance have on the dynamic target estimation performance, and what is the optimum allocation of RE and TDD gaps? The second results in an estimation problem from sparse data. Sparse allocation of REs and TDD gaps renders simple 2D-FFT target estimators useless since replacing the missing samples in the frequency domain just with zeros would distort the desired shape of the Ambiguity Function (AF) of the correlation reference waveform by disturbing sidelobes.
@@
The flexibility of 5G NR already allows for great waveform adaptability and optimum coordination of ISAC radio access. In addition to the physical OFDM parameters defined by the numerology, it is about the time-frequency distribution of the REs and TDD gaps in the radio frame, which is crucial for resource allocation. This enables the 5G communication system to accommodate a variety of applications with different bandwidth, latency, and reliability needs. We will observe a similar influence of time-frequency resource distribution for radar sensing quality control. ISAC radar sensing quality parameters are related to sensitivity of target detection, target resolution in range and Doppler, robustness against misdetection in clutter, target classification, and tracking performance in dynamic scenarios
@@
Here we will discuss some basic influential factors of MS access schemes and MS RE distribution. Figure 11 depicts the same three-node scenario as in Figure 10, however with two transmitters and one receiver. The two transmitters share the same OFDMA frame. We will recognize that sharing these resources also means a certain division of resolution performance for range and/or Doppler. The optimum sharing of OFDMA resources depends on the geometry of the sensor cluster relative to the target. In the left case, the direction of the target’s velocity vector is tangential to the Tx1 / Rx ellipse, while the right case aligns with the Tx2 / Rx ellipse. Provided the radio nodes are stationary, the observed Doppler shift would be zero for the Tx1 / Rx link in the left case and zero for the Tx2 / Rx in the right case. At the same time, in the corresponding other cases, the observed Doppler shift would be higher, although not necessarily maximum. The maximum Doppler shift would be observed when the velocity vector is perpendicular to the ellipses. This situation can be exploited at the transmitters for resource block allocation. Figure 12 depicts two examples of slow time sampling by OFDMA resource blocks with four transmitters. The left case describes uniform, repetitive sampling at the same rate for all four Tx. We see that shared Time-division Multiple Access (TDMA) access to the medium reduces the sampling rate in slow time and thus the maximum Doppler shift that can be estimated unambiguously. The included timing offset can be compensated for as long as the Nyquist sampling rate is maintained. The right part of Figure 12 explains how the TDMA scheme can be adjusted to the target dynamics. While Tx3 illuminates the target still with the same sampling rate in slow time, the unambiguously resolved Doppler bandwidth has reduced for Tx2 and Tx4. Tx1 can be used in two ways. If the two neighboring OFDM symbols are averaged, the effective overall Doppler sampling rate is kept, the Doppler bandwidth is reduced only a bit, and the sensitivity of target detection is increased. If the consecutive pairs of OFDM symbols undergo the FFT, the SNR gain by coherent averaging still applies, but we will also have some possibility to increase the resolvable Doppler frequency. The sensitivity of target detection can also be influenced by the radar integration time, i.e., the length of the slow time observation window, which can extend over several coherently processed OFDM frames. The limit is given by range migration, if the target moves from one range cell to the next one (in the range-Doppler domain).
@@
Figure 13 explains the FDMA split within the OFDM frame. The left part shows a uniform bandwidth allocation with compact bands that enable uniform range resolution for all four sensor links. In the right part, the same total bandwidth is allocated per transmitter, but the allocation is not consecutive, except for Tx4. A closer look shows that the range resolution capability for Tx1 is highest. However, because of the fragmented (”sparse”) frequency allocation, straightforward FFT estimation would not produce viable estimates because of grating lobes in the autocorrelation function. We would need a more sophisticated estimation procedure than just FFT that can cope with the sparse allocation of frequency resources as described in more detail in the next subsection.
@@
Some further comments are necessary. The time-frequency allocation described here as an example does not consider the time-frequency selectivity of the sensing channel. Especially the frequency selectivity because of multipath can require a more specific allocation of REs. Moreover, adaptive RE scheduling for joint optimization of sensing and communication quality of service would be necessary. Another interesting observation comes from the fact that for localization, we would need multiple measurements, e.g., at least 3 Tx / Rx pairs for 3D location and speed estimation. This means that the optimal distribution of resource blocks in the OFDM frame depends on the sensor geometry relative to the target and is mutually dependent. Moreover, for multiple targets, it depends on the difference of their dynamic target state vectors in its multiple dimensions. Therefore, the optimal allocation of resources in the OFDM frame leads to a joint optimization problem depending on the positions of all sensors used and the dynamic behavior of all relevant targets, given the available resources.
@@
Figure 11: Orthogonal MS access with Tx1 and Tx2 sharing the same OFDMA frame. In both cases, the position of the target is the same, but the direction of its movement is different.
@@
The discussion of MS-MIMO access in ISAC and any comparison with MU MIMO communication would be incomplete without considering Joint Transmission (JT), where the same illumination signal is transmitted by multiple distributed transmitters. Already throughout research towards 4G, Cooperative Multi-Point (CoMP) access, including JT, was discussed, especially for enhanced quality of service at cell edges. Although successfully demonstrated [ref33], it seems that CoMP has not been widely used up to now. However, the situation may develop differently for ISAC because of its inherent multilink requirements as explained above. JT for MS-MIMO ISAC definitely deserves attention in research. There seem to be remarkable differences to JT in communications. We have to make sure that focusing only involves the target and not the clutter. This could be achieved by background subtraction, e.g., if the target is moving and the clutter is static. However, due to the statistical reflectivity of typical extended targets, which leads to decorrelated radar returns as described above, coherent focusing to certain reflection points would only be possible for single Tx / Rx pairs. Therefore, although coherent JT promises tremendous spatial resolution that could resolve details from extended targets down to single reflection points, non-coherent focusing seems to be a more reasonable goal. This would lead to increased reflected energy off the target without increasing spatial resolution. At the same time, we get a diversity gain because of the different directions of the waves impinging on the target [ref34]. Thus, non-coherent JT ISAC seems to keep the statistical advantage of distributed MIMO radar, but the geometrical advantage may get lost since the multistatic Tx / Rx geometry cannot be traced back if the non-orthogonal waveforms cannot be decomposed at the receiver. However, since the Channel State Information (CSI) pilots are orthogonal, simultaneous estimation of the joint target channels is possible, which would allow adjusting the transmit signal delay in a subsequent step for non-coherent superposition as indicated in Figure 14.
@@ V-C Recovery of the Correlation Reference Signal
Excess time of arrival estimation, also jointly together with Doppler, can be based on specific pilot signals or on using the data payload as correlation reference to compare LOS path and BTP time of flight. Although pilots and other reference signals are designed for estimation of CSI in communications or UE positioning, they may not be enough for target state vector estimation in high dynamic range radar sensing. One reason for this is their sparse structure, which is sufficient for equalization, but interpolation errors in the frequency domain and the related distortion of the AF may not be tolerated for radar sensing. Moreover, since CSI pilots use only a fraction of the transmit signal power, the achievable correlation gain is limited. On the other hand, asking for more pilot signals is counterproductive, since this would reduce communication efficiency too much. Even extensive use of Position Reference Signal (PRS) would have the same effect. In contrast to this, CPCL as proposed in [ref1] and introduced already here in Section II, relies on the full data payload for precise target estimation and tracking. Note that radar localization has a systematic disadvantage compared to 3GPP UE device positioning since it relies on passive backscattering. This means that the sensing link budget follows the radar equation and not the Friis equation. Furthermore, the target cannot be identified by some user ID. Instead, we need to use much more sophisticated means to detect, recognize, and separate multiple targets. Using the data payload as correlation reference offers not only the possibility to use the total transmitted power (maximizing the SNR) and system bandwidth (maximizing resolution). It also allows the correlation reference waveform to be adjusted and modified, enabling unprecedented approaches to adaptive radar functions. However, the transmitted payload is not known a priori at the receiver (unlike the pilots) and has to be estimated at the receiver before performing correlation. While conventional passive radar (PCL) uses a dedicated, directional receiver channel to obtain a precise, multipath-free copy of the transmitted signal via LOS, ISAC’s own communication function offers many more options for recovering the transmitted signal without a separate Rx channel. However, we have to be careful in defining what we need as a reference and which recovery approach we apply. We will discuss different approaches and how they build upon the standard OFDM transmission approach.
@@
In addition to Tx / Rx frame and carrier synchronization and cyclic prefix removal, an OFDM receiver performs channel equalization in the frequency domain, which provides a symbol-by-symbol estimate of the transmitted OFDM signal based on multiplication by the inverse channel frequency response estimated at the CSI pilot positions and then interpolated to be applied to all carriers. This procedure is followed by decision making out of a finite modulation alphabet and by error correcting channel coding. As a result, the transmitted information can be decoded at the receiver with very low bit error probability. Therefore, we might come up with the idea to use that information for recovering the transmitted waveform for ISAC correlation by remodulation. However, the requirements for reliable and high dynamic range radar target detection and estimation are a bit different from communications. While in communication, we detect coded information per carrier; in radar, we perform a matched filter correlation using the whole OFDM symbol to identify the target range. The OFDM symbol is treated as an extended spread spectrum radar pulse that is compressed by correlation, which results in a rather high correlation gain. Moreover, we take further advantage of the coherent processing of many subsequent OFDM symbols. The result should clearly indicate the target-related propagation parameters delay and Doppler as described in the next subsection. However, the target may be associated only with a very small part of the received energy, perhaps in the neighborhood of strong clutter peaks and hidden by their correlation sidelobes. The dynamic range is therefore generally higher for ISAC than for communication. High detection probability and good resolution of target parameters in a continuous parameter space depend more on precise knowledge of the illumination waveform and its coherence over time than on its information content in terms of bits and bytes. However, RF impairments at the transmitter can affect the waveform and, hence, the shape of its correlation function. For instance, since with OFDM the power amplifier is often operated close to saturation, non-linear distortions can occur [ref37]. Also, linear distortions (complex frequency response, IQ imbalance, etc.) of the modulator, transmit filters, and antennas can prove to be disruptive. For communications, those effects are normally leveled out by the inherent error correction mechanisms. This may not be true for radar, as the detection problem is different. Targets often appear as small peaks in environments with strong clutter. They only become apparent when overlapping system responses can be resolved through iterative subtraction using a model-based estimator, as described in the next subsection. Therefore, precise knowledge of the shape of the transmitted waveform, as well as its correlation function (including the sidelobes), is of the utmost importance. This includes all distortions of the RF chain, including nonlinear ones that appear after digital reconstruction. Thus, it is preferable to rely entirely on the recovery of the correlation reference received over the air for target parameter estimation.
@@
Inverse filtering is the most important step in correlation reference recovery because it equalizes the frequency response, removes associated multipath, and provides a clean copy of the transmitted signal. However, the usual inverse filter operation in the frequency domain has serious pitfalls. One relates to REs (resp. carriers), which are loaded with zero power. First of all, these empty REs should not be included in the inverse filter operation. In the fully connected CPCL mode, the information about empty REs can be directly taken from the DeModulation Reference Signal (dmrs). In the more general case of the reduced CPCL mode, empty REs in the received reference signal are detected at the receiver based on a null hypothesis test [ref38]. Another related issue when recovering the transmit signal at the receiver side is the non-uniform (instantaneous) carrier magnitude. A simple inverse filter equalizes all received carriers to unit gain of the channel transfer function. This would lead to a non-uniform noise distribution of the carriers used for target parameter estimation. Since the CSI pilots are usually provided with more power and more robust modulation than the data carriers, these estimates would have a lower variance. On the other hand, multilevel quadrature amplitude modulation (QAM) of the carriers would lead to noise amplification at the carrier positions that are loaded according to the modulation alphabet by a lower magnitude of the constellation diagram. These effects have to be taken into account for optimum target parameter estimation as discussed in the next subsection.
@@
The quality of the reconstructed transmit signal can be further enhanced by a two-step iterative procedure. In a first step, we perform regular equalization of the received REs based on CSI pilots. The result is used for a first inverse filter step as described above. While these estimated payload data are processed for communication data reception, we use them to perform another channel equalization that should now be more accurate than the regular equalization since it uses the full transmit power and should be less affected by interpolation errors, which can arise because of the sparse structure of CSI pilots. With the now enhanced reference signal, we can calculate the final channel impulse response, which is used for radar target detection and estimation. The advantage of this approach compared to the alternative based on transmit symbol decision, decoding, and remodulation at the receiver is as follows. All linear and non-linear distortions at the transmitter are now included in the correlation reference, which assures a high dynamic range estimate. Even the time reference is kept, and Tx2Rx synchronization is automatically established since the critical eToF and eDoppler are effectively calculated as differences between LOS and BTP from the same measured CIR (see discussion on synchronization issues later on). We refer to this approach as Turbo CPCL because it is a nested iterative approach that takes advantage of the processing gain in the inner estimation loop using the CSI pilots. In terms of estimation accuracy, we can also understand the outer loop as a refinement of the first estimate.